# Exploration of Masked and Causal Language Modelling for Text Generation

### v1

@@ III-D3 Authorship Verification Task
Authorship verification is the task of analysing the linguistic patterns of two or more texts to determine whether the same author wrote them. Initially, experts performed this analysis by considering spelling mistakes, grammatical inconsistencies and other linguistic features. However, more recently, machine learning algorithms have been trained to accomplish the same. In this task, we collected all the pairs of texts that our fine-tuned model originally predicted belonged to the same author 4. We then altered one of the texts in each pair, and we tried to determine if the prediction of the model about the authors being the same remained unchanged despite the corruption.
@@
As Figure 4 shows, for this task first, we give 2 a pair of texts to the classification model. If the model outputs "Yes", we then corrupt Text 2 and reconstruct it using our generative model. Finally, we classify the pair again and count the number of times the model correctly outputs "Yes" again. We focused only on when the model outputs "Yes" initially because it is harder for two texts to come from the same author than the opposite.
The third downstream task does not directly involve training models and instead it essentially aims to determine whether the generated text was of high quality enough to "fool" the prediction model into thinking that the text still comes from the same author by maintaining a similar writing style to the original text.
Author Verification Model: Flan-T5-XL Fine-tuned LAnguage Net T5 XL (FLAN-T5-XL) [31] is an extension of the T5 model designed for tasks requiring a deep understanding of language context. It is trained on the "text-to-text" framework (see Section II-A2). The model was developed to address a range of text-based tasks, including translation, summarisation, question answering, and text generation. In this context, the "XL" stands for the size of the model, which is 3 billion parameters in this case. The model has already been trained on the Authorship Verification task.
@@ IV Results and Discussion
In the following section, we first show the results of the quantitative, qualitative, and downstream evaluations. We then discuss the results and their possible implications. Due to computational constraints when generating text with MLM (as explained in Section V-A), the synthetic datasets were between 130 and 200 samples in size.
@@ IV-A Quantitative Evaluation
After computing the scores for the quantitative evaluation, we plotted them in the following figures. The evaluation metrics we used compare the generations obtained from corrupted text and the original texts before they were corrupted. The scores of all of the masking approaches (introduced in Section III-C3) we used are present in the figures; scores from random masking are shown as the masking ratio increases, while scores obtained by the other approaches are added to the graph as flat lines. See Appendix LABEL:app:quantitative_results_graphs for the complete collection of figures for each dataset. The MLM models are RoBERTa, BiomedNLP-PubMedBERT, and BERT, whilst the CLM models are T5, SciFive, and BART.
@@ IV-A1 Medical Dataset
As shown in Figure 5, the MLM model generations (Roberta-large & BiomedNLP-PubMedBERT) align more closely with the original text and, therefore, achieve higher scores. It is clear that MLM models are more sensitive to changes in the masking ratio, and they get better scores when the ratio is lower. They also show very low scores when the ratio is 100%. With regard to CLM models, they show low scores across all generation approaches; in this case, a low BLEU score shows low precision. The punctuation and stopwords masking are also relatively good for the MLM models showing mid-range scores, while in CLM, they are at the bottom of the graph.
@@
Figure 6 shows that while still worse than the MLM models, the ROUGE scores of the CLM models are relatively better than the BLEU scores, meaning that at least some of the words that appear in their generations also appear in the reference texts (recall).
@@
As shown in Figure 7, the METEOR evaluation, a recall-oriented metric, confirms our hypothesis above, showing relatively good scores for the CLM models compared to the BLEU scores. However, it is clear that CLM models include many tokens in their generations that do not appear in the reference texts compared to the MLM models.
@@
In Figure 8, the BERTScore shows that notwithstanding the specific medical language, all models can generate samples that are semantically similar to the reference texts. This shows good adaptability to the task-specific domain for all models, with MLM models showing slightly better performance.
@@ IV-A2 Movie Dataset
Again, as expected, Figure 9 less masking leads to higher scores. When generating samples from the movie dataset, the MLM model still achieves higher scores. While t5-large’s score is as low as with the medical dataset, bart-large improved significantly, achieving good scores overall. In addition, the model seems more sensitive to the level of context, meaning that the model can use it better and produce informed generations. Neither the punctuation nor stopwords approaches provide enough context for the model to generate good-quality samples, with their scores being at the bottom of the graph.
IV-A3 Author Dataset
@@
As with the other two datasets, Figure 10 shows that MLM models perform better than the CLM models; this time, bart-large performs well in the variable rations compared to the previous two datasets, while t5-large is not able to produce accurate generations.
@@ IV-B Human Evaluation on Sampled Data
For this task, we picked the first 5 samples generated by the models when using a random 50% masking ratio.
@@ IV-B1 Medical Generations
MLM Generations These samples were generated by Roberta large.
@@
As shown in Table III, MLM generations are high in quality. They show few hallucinations, a high level of coherence in the texts, and a good length that matches the expected length for discharge summaries. In addition, these texts show an appropriate structure for medical discharge summaries because they include relevant sections such as patient instructions, medication changes, and follow-up recommendations. What is also remarkable is that the model demonstrates a good grasp of medical terms, such as "angioedema," "pneumonia," and "endoscopy." The texts are generally coherent; however, information is sometimes introduced out of nowhere. For example, in the fourth summary, "mm nodular area on the chest CT" is disconnected from the rest of the summary.
CLM Generations These samples were generated by T5 large.
be at follow up please call with any or concerns/fax cardiac surgery office any or concerns/fax cardiac surgery office any or concerns telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone.’, 1: ’take any new. Please take any new of regular medications take any new as ordered. Please of regular medications take any new. Please of regular medications take any new. Please of regular medications take any new. Please of regular medications take any new of regular medications take any new. Please. Please to to. Please to chest pain to to chest to chest to chest to to……. take….
@@
In Table IV, the texts show many repetitions; multiple phrases are repeated for almost the entire length of the texts. For example, in the first generated text, the phrase "answering service contact on hours telephone" is repeated multiple times. First of all, the generated texts lack the structure and organization that proper medical discharge summaries would require; second, most of the information appears randomly generated and without a clear flow. This all together makes it very difficult to understand the key instructions given by the summary. Some texts contain incomplete sentences, such as "take any new. Please take any new or regular medications as ordered. Please, of regular medications, take any new." Over longer sentences, it is clear that the CLM model cannot maintain any coherence in the texts; indeed, while the texts include medical terms and phrases, they are often used in a disjointed way and lack context. For instance, "systolic. blood systolic. blood systolic. blood systolic. blood systolic. blood primary care doctor to hospital with a reaction called angioedema." is a clear hallucination. There are many instances of irrelevant information being generated, such as "if it becomes black or blood in stool, go to the hospital." which appears in the summary without appropriate context.
@@ IV-B2 Movie Generations
MLM Generations These samples were generated by Roberta large.
1. The film opens with a montage of scenes of daily life in Australia during the apocalypse: a rural school in the desert, a busy street in a small town, a traffic jam in the city, all being affected by the adverse weather conditions that suddenly appear. Only the local Aboriginal people seem to recognize the changes. The film’s first scene is one of these "freak occurrences" in Sydney, where it occurs among a group of men in a pub, all affected by the mysterious effects of the weather. Arriving at the court of public safety, the case is ruled not guilty; and the men are given the opportunity to use the public legal aid system, where Burton is the only legal counsel procured for their defense. The circumstances under which he was contacted and retained are unusual, in that his law practice is corporate, and not criminal defense. The film focuses on the case, and the professional and personal lives of the unravel involved, including bizarre dreams that Burton relates to the court. Interacting interviewing one of the accused (David Gulpilil), he is also able to uncover a strange ritual that has taken place in his home, and connect his rituals with the murder case, which turns out to have been an Aboriginal ritual performed by curse same shaman that the victim believed. Learning more about Aboriginal practices and the idea of Dreamtime as a parallel world of existence, Burton concludes that this strange ritual bodes ill for the apocalypse. The film ends with a confrontation between Burton and the victim’s shaman at a subterranean burial site. Overcoming the shaman and Burton reach the surface, buried in a tunnel are various tribal relics. After emerging from the tunnel, he sits on the beach and stares out at the horizon. The screen is filled with the shot of a crashing ocean wave, though it is unclear whether we are witnessing the end of the world’s final, apocalyptic storm.
@@
Table V shows that the generated texts have different levels of coherence and grammatical correctness. Overall, the summaries convey the main storylines smoothly and introduce the central characters at the right time; however, the narration is disrupted multiple times because of the hallucinations introduced by the model.
Many of the transitions in the summaries happened all of a sudden, which makes it harder to follow the plots. The second summary, for example, jumps from the childhood of the protagonist to his adult life. In addition, the generated text sometimes includes irrelevant details that draw attention away from the main storyline, such as the "Glock" and the slow movement of the desk in the fifth summary, which seem totally out of context. The model is sometimes limited in its ability to maintain long-term dependencies between characters in the plots; for example, in the fourth plot, the summary contradicts itself multiple times about the actions of Duke and Meta.
Nonetheless, the generations are mostly understandable at least. The main characters, settings, and conflicts are all present in the summaries. They are also easy to read, have clear language, and are similar to the language used in the original training set. The summaries are very different in style and tone, showing that the models can adapt to different genres.
CLM Generations These samples were generated by Bart large.
With montage scenes of daily Australia 1970s: desert main of a jam in the, being unusually adverse weather that appear only the aboriginal to recognize the significance of weather. The Australian legal aid, (Chamberlain) defence the circumstances by which, in that taxation the and personal lives unravel. He is plagued by bizarre dreams begins to connection to one of the accused (David Gulpilil) also feels connected increasingly strange phenomena city. His intensify along with the case, which he comes an aboriginal killing curse in which the victim about aboriginal and the as a parallel world of existence, Burton comes to weather bodes the end of the world. Overcoming the climaxes confrontation the lawyer and tribe’ subterranean sacred site. After the sewer, he on beach and stares. Screen filled by the shot of a ocean wave unclear if he is witnessing or in’ final, he stares at the ocean. The film ends with one of these "freak" Sydney an occurs among a of in the mysterious of the mysterious death of one of at coroner, death ruled homicide; men. Burton escapes to in the tunnel various relics. Overcoming, he escapes to the beach, where he stares to the ocean and stares at it.
@@
Table VI shows that overall, the quality of these writings is far below that of the MLM generations. The primary issue is the lack of coherence in specific passages as the text occasionally struggles to maintain a clear narrative flow, with abrupt transitions and unclear connections between each sequence. Most passages contain vague statements that leave room for interpretation; they seem more like a collection of random sentences lacking logical connections between them. Another critical problem is that the texts contain grammatical errors, such as incorrect verb tenses, subject-verb agreement issues, and awkward sentence structures. Additionally, the text sometimes lacks clarity and precision in its language, with the occasional appearance of random characters in the text.
@@ IV-B3 Author Generations
MLM Generations These samples were generated by Roberta large.
’hi, thank you so much for clarifying this, im so sorry for the inconvenience i am happy to take all the shifts for tomorrow, again thank you for your help Hi, good afternoon, i have been unable to upload my powerpoint presentation as an attachment as the file is not working i have attempted to upload it as an onedrive, however i am experiencing issues could i please upload the slides from my memory stick if any further attempts fail i will continue to continue to attempt to upload if you could let me know if this is working i would be very grateful kindest regards, Hi, i have estimated as an end date for tomorrow hope this is okay truly, Hi, i have attached pictures of my presentation, as have many others, good afternoon, i hope your well im sorry for the delayed reply, i was unable to hear back from you guys about my dbs, which has now been received i have scanned my dbs but they are much bigger than a4x5 printer normally scans so they will not show all the information on each side i was able to print out all the documents on hard drive i have attached the pdf containing the information required per the interview, there are also 2 pdfs containing my CV and email address thank you for your help, Hi, my lovelies, i hope youre all well i am so happy to inform you i have secured the role with i would not have got through the interview without all your support and help, i cant thank you enough for your help with my placement i hope you guys will be able to answer if i am missing anything in any of my questions, i can tell you that i will be completing my placement for 100 days, rather than a year anything which is very useful to carry around day i am currently searching for a diary what is training and is it something i should bringwear me as i havent been provided with a location yet, will routes also be required to attend training, or will it depend on the number of days i will be on a placement is there anything you can help me with before starting, any information i can get is much more beneficial to me i will be contacted via your website, once again thank you for your continuous,’
@@
Table VII shows that one significant issue throughout the texts is the lack of proper grammar, punctuation, and sentence structure. The model struggles to maintain coherence and clarity, often producing fragmented and poorly organized ideas. In addition the texts are full of repetitive phrases, lack of smooth transitions between ideas. However, the model does have the ability to capture different writing styles, such as casual conversations or more formal tones. Overall, the texts are understandable and do not show any significant errors.
Hi thank you so clarifying that im will bring the for once thank you, good afternoon, i am unable to my powerpoint presentation as attachment as the is too big to it as onedrive i am experiencing some issues with my memory, could i the presentation memory all other attempts fail i will however to you this is i be very grateful kindest regards, date placement i hope yours, hi have attached dbs thanks, Hi , I have attached my dbs Thank you for your support, Hi, I hope youre well, I have a queries the placement hope will be answer if thats ok to of my questions can tell you that i be for days over a will to carry day to day i am currently what training like should i been with location yet so routes would be helpful recommendations the prices on number days i will completing you guys before placement any you think relevant for to know i contacted this again thank you for my support, thank you enough Good afternoon, My name is and I am a student at , I am on placement for with I am so pleased to the role with i not gone through the application process as i have not received any support, Thank you enough, Hello, your have attached the dbs to my presentation Thanks, Good morning, my university about form has now i have as bigger than a4 only a4 show all each am to documents as copies 4 the required per there are also 2 attached containing address you thanks, hi, i hope Youre well im delayed i have been contacted by the university about my form, my university has now informed me that my form has been sent to for me to have a look at it thanks,
@@
CLM Generations These samples were generated by Bart large. In Table VIII, the texts cover various topics and styles, including personal communication, academic writing, and conversational transcripts. However, the overall quality of the generated content is poor and lacks coherence, proper grammar, and punctuation. As with the movie synopses, the model struggles to maintain clarity, generating fragmented sentences. The information presented is often confusing and contains random or irrelevant details. Overall, the models cannot maintain a clear focus and coherent flow throughout each text.

### v2

@@ III-D3 Authorship Verification Task
Authorship verification involves the examination of linguistic patterns in two or more texts to verify whether they were authored by the same individual. Traditionally, experts conducted this analysis by considering spelling errors, grammatical inconsistencies, and other linguistic features. In recent times, however, machine learning algorithms have been trained to perform this task. As shown in Figure 5, we compiled all pairs of texts that our fine-tuned model initially identified as having the same author. Subsequently, we modified one text in each pair and assessed whether the model’s prediction regarding authorship remained consistent despite these alterations.
@@
Note that this third downstream task does not directly involve training models, but instead essentially aims to determine whether the generated text was of high quality enough to fool the predictive model into thinking that the text still comes from the same author by maintaining a similar writing style to the original text.
@@ IV Results and Discussion
In the following section, we first show the results of the quantitative, qualitative, and downstream evaluations. We then discuss the results and their possible implications. Due to computational constraints when generating texts with MLM (as described in Section V-A), the synthetic datasets were restricted between 130 and 200 samples.
Figure 6: Evaluation Pipeline.
@@ IV-A Quantitative Evaluation
After computing the scores for the quantitative evaluation, we visualized them in the subsequent figures. The evaluation metrics compare the generated texts derived from corrupted texts with the original texts before corruption. The figures include scores from all masking approaches discussed in Section III-C3. For random masking, scores are displayed as the masking ratio increases, while scores from other approaches are depicted as flat lines. The MLM models used are RoBERTa, BiomedNLP-PubMedBERT, and BERT, whereas the CLM models include T5, SciFive, and BART.
@@ IV-A1 Medical Dataset
As illustrated in Figure 10, the generations from the MLM models (RoBERTa-large and BiomedNLP-PubMedBERT) align more closely with the original text, thereby achieving higher scores. These models are more sensitive to changes in the masking ratio, performing better when the ratio is lower, and exhibiting very low scores when the ratio reaches 100%. In contrast, the CLM models display low scores across all generation approaches, indicating low precision as reflected by the low BLEU scores. For the MLM models, masking punctuation and stopwords results in mid-range scores, whereas for the CLM models, these approaches yield scores at the bottom of the graph.
Figure 10 indicates that, although still inferior to the MLM models, the ROUGE scores of the CLM models are relatively better than their BLEU scores. This suggests that at least some of the words generated by the CLM models also appear in the reference texts, indicating a higher recall.
Figure 10 supports this observation through the METEOR evaluation, a recall-oriented metric, which shows relatively good scores for the CLM models compared to their BLEU scores. However, CLM models generate many tokens that are not present in the reference texts, unlike MLM models.
In Figure 10, the BERTScore reveals that despite the specific medical language, all models can generate samples semantically similar to the reference texts. This demonstrates good adaptability to the task-specific domain for all models, with MLM models exhibiting slightly better performance.
@@ IV-A2 Movie Dataset
As observed in Figure 12, less masking results in higher scores. When generating samples from the movie dataset, the MLM model continues to achieve higher scores. Although T5-large’s scores remain as low as those with the medical dataset, BART-large shows significant improvement, achieving good overall scores. Additionally, the model appears more sensitive to the level of context, suggesting it can better utilize context to produce informed generations. Neither the punctuation nor stopwords masking approaches provide sufficient context for generating high-quality samples, as evidenced by their scores being at the bottom of the graph.
IV-A3 Author Dataset
Once again, Figure 12 demonstrates that MLM models outperform CLM models. Bart-large performs well through the variable ratios compared to the previous two datasets, while t5-large cannot produce accurate generations.
@@ IV-B Human Evaluation on Sampled Data
In this subsection, we pick the first sample generated by each model using a random 50% masking ratio.
@@ IV-B1 Medical Generations
MLM Generations. As shown in Table III, MLM sequences generated by Roberta-large are indeed of high quality. Although they show a few hallucinations, we observe a high level of coherence in the texts and sequences that match the expected length for discharge summaries. In addition, these texts show an appropriate structure for medical discharge summaries as they include relevant sections such as patient instructions, medication changes, and follow-up recommendations. What is also remarkable is that the model demonstrates a good grasp of medical terms, such as "angioedema," "pneumonia," and "endoscopy". The texts are generally coherent; however, information is sometimes introduced out of nowhere. For example, in the fourth summary, "mm nodular area on the chest CT" is disconnected from the rest of the summary.
@@
CLM Generations.
be at follow up please call with any or concerns/fax cardiac surgery office any or concerns/fax cardiac surgery office any or concerns telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone.’, 1: ’take any new. Please take any new of regular medications take any new as ordered. Please of regular medications take any new. Please of regular medications take any new. Please of regular medications take any new. Please of regular medications take any new of regular medications take any new…
@@
In Table IV, the texts generated by T5-large show many repetitions where multiple phrases are repeated over the entire length of the sequence. Note, for example, the phrase "answering service contact on hours telephone" that is repeated twelve times over the entire sequence. First of all, the generated texts lack the structure and organization that proper medical discharge summaries would require. Second, most of the information appears randomly generated and without a clear flow. This all together makes it very difficult to understand the key instructions given by the summary. In addition, some texts contain incomplete sentences, such as "take any new. Please take any new or regular medications as ordered. Please, of regular medications, take any new.". Over longer sentences, it is clear that the CLM model cannot maintain any coherence in its generated texts; indeed, while the texts include medical terms and phrases, they are often used in a disjointed way and lack context. For instance, "blood systolic. blood systolic. blood primary care doctor to hospital with a reaction called angioedema." is a clear example of hallucination. Finally, we can also observe many instances of irrelevant information being generated, such as "if it becomes black or blood in stool, go to the hospital.", which appears in the summary without appropriate context.
@@ IV-B2 Movie Generations
MLM Generations. Table V shows that the texts generated by Roberta-large have different levels of coherence and grammatical correctness. Overall, the summaries convey the main storylines smoothly and introduce the central characters at the right time; however, the narration is disrupted multiple times because of the hallucinations introduced by the model.
Many of the transitions in the summaries happened all of a sudden, which makes it harder to follow the plots. The second summary, for example, jumps from the childhood of the protagonist to his adult life. In addition, the generated text sometimes includes irrelevant details that draw attention away from the main storyline, such as the "Glock" and the slow movement of the desk in the fifth summary, which is totally out of context. The model is sometimes limited in its ability to maintain long-term dependencies between characters in the plots; for example, in the fourth plot, the summary contradicts itself multiple times about the actions of Duke and Meta.
Nonetheless, the generations are mostly understandable. The main characters, settings, and conflicts are all present in the summaries. They are also easy to read, have clear language, and use languages similar to the ones used in the original training dataset. Finally, the summaries are very different in style and tone, showing that the models can adapt to different genres.
The film opens with a montage of scenes of daily life in Australia during the apocalypse: a rural school in the desert, a busy street in a small town, a traffic jam in the city, all being affected by the adverse weather conditions that suddenly appear. Only the local Aboriginal people seem to recognize the changes. The film’s first scene is one of these "freak occurrences" in Sydney, where it occurs among a group of men in a pub, all affected by the mysterious effects of the weather. Arriving at the court of public safety, the case is ruled not guilty; and the men are given the opportunity to use the public legal aid system, where Burton is the only legal counsel procured for their defense. The circumstances under which he was contacted and retained are unusual, in that his law practice is corporate, and not criminal defense. The film focuses on the case, and the professional and personal lives of the unravel involved, including bizarre dreams that Burton relates to the court. Interacting interviewing one of the accused (David Gulpilil), he is also able to uncover a strange ritual that has taken place in his home, and connect his rituals with the murder case, which turns out to have been an Aboriginal ritual performed by curse same shaman that the victim believed. Learning more about Aboriginal practices and the idea of Dreamtime as a parallel world of existence, Burton concludes that this strange ritual bodes ill for the apocalypse. The film ends with a confrontation between Burton and the victim’s shaman at a subterranean burial site.
@@
CLM Generations.
With montage scenes of daily Australia 1970s: desert main of a jam in the, being unusually adverse weather that appear only the aboriginal to recognize the significance of weather. The Australian legal aid, (Chamberlain) defence the circumstances by which, in that taxation the and personal lives unravel. He is plagued by bizarre dreams begins to connection to one of the accused (David Gulpilil) also feels connected increasingly strange phenomena city. His intensify along with the case, which he comes an aboriginal killing curse in which the victim about aboriginal and the as a parallel world of existence, Burton comes to weather bodes the end of the world. Overcoming the climaxes confrontation the lawyer and tribe’ subterranean sacred site. After the sewer, he on beach and stares. Screen filled by the shot of a ocean wave unclear if he is witnessing or in’ final, he stares at the ocean. The film ends with one of these "freak" Sydney an occurs among a of in the mysterious of the mysterious death of one of at coroner, death ruled homicide; men. Burton escapes to in the tunnel various relics.
@@
Table VI shows that overall, the quality of the writings produced by Bart-large is far below that of the MLM generations. The primary issue is the lack of coherence in specific passages as the text occasionally struggles to maintain a clear narrative flow, with abrupt transitions and unclear connections between each sequence. Most passages contain vague statements that leave room for interpretation and seem more like a collection of random sentences lacking logical connections between them. Another critical issue is that some sequences contain grammatical errors, such as incorrect verb tenses, subject-verb agreement, and awkward sentence structures. Additionally, the texts sometimes lack clarity and precision in their language, with some occasional appearance of random characters.
@@ IV-B3 Author Generations
MLM Generations.
’hi, thank you so much for clarifying this, im so sorry for the inconvenience i am happy to take all the shifts for tomorrow, again thank you for your help Hi, good afternoon, i have been unable to upload my powerpoint presentation as an attachment as the file is not working i have attempted to upload it as an onedrive, however i am experiencing issues could i please upload the slides from my memory stick if any further attempts fail i will continue to continue to attempt to upload if you could let me know if this is working i would be very grateful kindest regards, Hi, i have estimated as an end date for tomorrow hope this is okay truly, Hi, i have attached pictures of my presentation, as have many others, good afternoon, i hope your well im sorry for the delayed reply, i was unable to hear back from you guys about my dbs, which has now been received i have scanned my dbs but they are much bigger than a4x5 printer normally scans so they will not show all the information on each side i was able to print out all the documents on hard drive i have attached the pdf containing the information required per the interview, there are also 2 pdfs containing my CV and email address thank you for your help, Hi, my lovelies, i hope youre all well i am so happy to inform you i have secured the role with i would not have got through the interview without all your support and help, i cant thank you enough for your help with my placement i hope you guys will be able to answer if i am missing anything in any of my questions, i can tell you that i will be completing my placement for 100 days, rather than a year anything which is very useful to carry around day i am currently searching for a diary what is training and is it something i should bringwear me as i havent been provided with a location yet,’
@@
Table VII shows that one significant issue within texts generated by Roberta-large is the lack of proper grammar, punctuation and sentence structure. The model struggles to maintain coherence and clarity, often producing fragmented and poorly organized ideas. In addition, the texts are mostly made up of repeated phrases, lacking smooth transitions between ideas. However, the model does have the ability to capture different writing styles, such as casual conversations or more formal tones. Overall, the generated texts are understandable and do not include significant errors.
Hi thank you so clarifying that im will bring the for once thank you, good afternoon, i am unable to my powerpoint presentation as attachment as the is too big to it as onedrive i am experiencing some issues with my memory, could i the presentation memory all other attempts fail i will however to you this is i be very grateful kindest regards, date placement i hope yours, hi have attached dbs thanks, Hi , I have attached my dbs Thank you for your support, Hi, I hope youre well, I have a queries the placement hope will be answer if thats ok to of my questions can tell you that i be for days over a will to carry day to day i am currently what training like should i been with location yet so routes would be helpful recommendations the prices on number days i will completing you guys before placement any you think relevant for to know i contacted this again thank you for my support, thank you enough Good afternoon, My name is and I am a student at , I am on placement for with I am so pleased to the role with i not gone through the application process as i have not received any support, Thank you enough, Hello, your have attached the dbs to my presentation Thanks, Good morning, my university about form has now i have as bigger than a4 only a4 show all each am to documents as copies 4 the required per there are also 2 attached containing address
@@
CLM Generations. Table VIII displays texts generated by Bart-large. As we can observe, the sequences cover various topics and styles, including personal communication, academic writing, and conversational transcripts. However, the overall quality of the generated content is poor and lacks coherence, proper grammar, and punctuation. As with the movie synopses, the model struggles to maintain clarity, generating fragmented sentences. The information presented is often confusing and contains random or irrelevant details. Overall, the models fail to maintain a core idea, demonstrating a lack of coherent flow throughout their generations.